Evaluate


int(9/4) (dx/✓x)


1
Expert's answer
2021-11-15T16:23:10-0500

Solution:

"I=\int_4^9\dfrac1{\sqrt x}dx
\\=\int_4^9 (x)^{-1/2}dx
\\=\dfrac{x^{1/2}}{1/2}|_4^9
\\=2\sqrt x|_4^9
\\=2(\sqrt 9-\sqrt 4)
\\=2(3-2)
\\=2(1)
\\=2"